Β have a mass of 8000g and the total area of my feet is 0.04 square meters




1
Expert's answer
2021-02-24T12:47:56-0500
"p=\frac{F}{A}\\p=\frac{mg}{A}\\p=\frac{8\cdot9.8}{0.04}=1960\ Pa"